OpenStudy (anonymous):

50 is what percent of 80

OpenStudy (anonymous):

\[\frac{50}{80}\times 100 \]percent

OpenStudy (anonymous):

\[=\frac{5}{8}\times 100=0.625 \times 100 = 62.5\]percent